The perimeter of a square is 13 feet, 8 inches. What is the length of one side of the square?... An please explain in details.
Answer by nerdybill(7384)

The perimeter of a square is 13 feet, 8 inches. What is the length of one side of the square?... An please explain in details.
.
12 inches per feet
so,
13 feet = 13*12 = 156 inches
.
13 feet, 8 inches = 156 + 8 = 164 inches
.
Since it is a square, we know that all four sides are the same.
Let x = length of one side of square
then
4x = 164
x = 164/4
x = 41 inches